# ScaleGNN: Towards Scalable Graph Neural Networks via Adaptive High-order Neighboring Feature Fusion

### v2

@@ 4. Methodology
In this section, we introduce ScaleGNN, an efficient end-to-end framework designed to address the scalability and performance challenges of GNNs. Our proposed ScaleGNN integrates several key innovations to handle issues such as over-smoothing, computational inefficiency, and redundant high-order information. The framework consists of three core components: (1) Adaptive High-order Feature Fusion, (2) Low-order Enhanced Feature Aggregation, and (3) High-order Redundant Feature Masking. These components are designed to work in synergy, allowing ScaleGNN to efficiently process large-scale graphs while maintaining effective information flow across different neighborhood orders. The outline of our ScaleGNN is shown in Fig. 1.
4.1. Adaptive High-order Feature Fusion
@@
Traditional GNNs are based on message passing through fixed adjacency matrices, where the adjacency structure remains constant throughout the layers. This approach can lead to over-smoothing, where node representations converge and lose their distinctiveness as the network depth increases. Furthermore, as nodes aggregate information from higher-order neighbors, the feature similarity across nodes can grow excessively, reducing the expressiveness of the model.
@@
To mitigate these issues, we propose a method for constructing distinct adjacency matrices for each neighborhood order. This allows us to isolate the relationships specific to each order, ensuring that each neighborhood level contributes uniquely to the node representation. We define the pure $i$-th order adjacency matrix as the difference between the $(i-1)$-th and $i$-th powers of the adjacency matrix:
(1) $\mathbf{A}_{i}=\mathbf{A}^{i}-\mathbf{A}^{i-1},\quad i=1,2,3,\dots,K,$
@@
where $\mathbf{A}^{i}$ is the $n$-th power of the adjacency matrix, which captures the connections up to $n$ hops, and $\mathbf{A}_{i}$ isolates the new relationships emerging specifically at the $i$-th order. Specifically, $\mathbf{A}_{1}$ is the adjacency matrix of the graph, $\mathbf{A}_{0}$ is the unit matrix.
@@
Incorporating information from multiple orders can provide a more comprehensive view of the graph structure. However, the challenge lies in integrating these multiple orders in a way that preserves their distinct contributions. To address this, we introduce a learnable weight vector $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\dots,\alpha_{K}]$, where each $\alpha_{i}$ represents the importance of the $i$-th order adjacency matrix in the final aggregation:
(2) $\tilde{\mathbf{A}}=\sum_{i=1}^{K}\alpha_{i}\mathbf{A}_{i},\quad\sum_{i=1}^{K}\alpha_{i}=1.$
@@
These weights are learned through back-propagation, ensuring that the model dynamically adjusts the contribution of each neighborhood order. The use of the softmax function to optimize these weights ensures that the contributions across the different orders remain balanced. Finally, we obtain the adaptive high-order fused feature matrix $\mathbf{H}\in\mathbb{R}^{n\times d}$ as follows:
(3) $\mathbf{H}=\sigma(\tilde{\mathbf{A}}\cdot\mathcal{X}\cdot\mathbf{W}),$
@@
where $d$ is the hidden embedding dimension, $\mathcal{X}\in\mathbb{R}^{n\times f}$ is the input feature matrix, and $\mathbf{W}$ is learnable weight matrix for the high-order features.
@@
We use the model obtained here as the base version named ScaleGNNb, which is relatively simple and efficient.
4.2. Low-order Enhanced Feature Aggregation
@@
In many GNN models, the integration of low-order (local) and high-order (global) features is done in a single, often static, step. However, this can lead to high-order features dominating over low-order ones, which might diminish the model’s ability to preserve fine-grained, local information. To address this, we propose to explicitly separate the low-order and high-order features before fusion.
First, we obtain the low-order feature matrix $\mathbf{H}_{\text{low}}$ and the high-order feature matrix $\mathbf{H}_{\text{high}}$:
@@
(4) $\mathbf{H}_{\text{low}}=\sigma(\mathbf{A}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{low}}),\quad\mathbf{H}_{\text{high}}=\sigma(\tilde{\mathbf{A}}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{high}}),$
@@
where $\mathbf{W}_{\text{low}}\in\mathbb{R}^{f\times d}$ and $\mathbf{W}_{\text{high}}\in\mathbb{R}^{f\times d}$ are both learnable weight matrices for the low-order and high-order features, respectively. The activation function $\sigma(\cdot)$ introduces non-linearity to the transformation.
@@
To ensure a balanced integration of local and global information, we introduce a learnable balancing factor $\beta$, which allows for dynamic adjustment of the contribution from each feature type. The final node representation $\mathbf{H}$ is computed as a weighted sum of the low-order and high-order features:
(5) $\mathbf{H}=\beta\cdot\mathbf{H}_{\text{low}}+(1-\beta)\cdot\mathbf{H}_{\text{high}},$
@@
where $\beta$ is optimized during training to ensure that the model can adaptively balance the trade-off between local and global information, depending on the specific task and graph structure.
@@
4.3. High-order Redundant Feature Masking
Higher-order neighbors introduce diverse, yet potentially redundant, information to the GNN. Without an effective masking mechanism, these neighbors can introduce noise, diluting the informative signal and impairing the model’s ability to generate useful node embeddings. To address this, we propose a novel method for selecting the most relevant and masking other redundant high-order neighbors based on their local contribution to node representations.
We introduce the Local Contribution Score (LCS) to assign importance scores to high-order neighbors. The LCS is designed to quantify the relevance of each neighbor by considering both structural similarity and feature alignment. Specifically, for a node $v_{i}$ and its $k$-th order neighbor $v_{j}$, the LCS score is computed as:
(6) $\text{LCS}(v,u,k)=\frac{\alpha_{k}(\mathbf{A}_{k})_{vu}\cdot(x_{v}^{\top}x_{u})}{\sqrt{d_{v}\cdot d_{u}}},$
where $(\mathbf{A}_{k})_{vu}$ is the strength of the connection between nodes $v$ and $u$ in the $k$-hop adjacency matrix, $x_{v}$ and $x_{u}$ are the feature vectors of nodes $v$ and $u$, respectively, and $d_{v}$ and $d_{u}$ represent the degrees of nodes $v$ and $u$, which normalize the influence of highly connected nodes. $\alpha_{k}$ is the learnable parameter that controls the contribution of the $k$-th hop’s adjacency matrix defined in Eq. 2.
The LCS allows us to mask redundant high-order neighbors and identify the most relevant neighbors by considering both local structure and feature similarity, enabling the model to selectively pass information from the most meaningful high-order neighbors. Once the LCS is computed for each neighbor, we select the top-$K$ neighbors with the highest LCS values for each node:
(7) $\mathbf{A}_{k}^{\text{filtered}}=\text{Top}_{K}(\mathbf{A}_{k},\text{LCS}(u,v,k)).$
This filtering process ensures that only the most informative high-order neighbors contribute to the node’s final representation, effectively reducing noise and improving the stability of the embeddings.
4.4. Optimization Objective
To ensure the effectiveness of our proposed ScaleGNN, we design an optimization framework that consists of three key loss components: the primary task loss, the local contribution score (LCS) regularization, and the mask constraint on high-order information masking mechanism. These losses work together to enhance model performance, suppress noise from uninformative high-order neighbors, and maintain computational efficiency.
@@ 4.4. Optimization Objective
(8) $\mathcal{L}_{\text{task}}=-\sum_{i\in\mathcal{V}}y_{i}\log\hat{y}_{i},$
where $y_{i}$ represents the ground truth label for node $i$, and $\hat{y}_{i}$ is the predicted probability distribution over classes.
To guide the model in selecting meaningful high-order neighbors, we introduce an LCS-based regularization term:
(9) $\mathcal{L}_{\text{LCS}}=\sum_{i=2}^{K}\sum_{(v,u)\in\mathcal{N}_{i}^{\text{selected}}}\left(1-\text{LCS}(v,u)\right)^{2}.$
where $\mathcal{N}_{i}^{\text{selected}}$ is the selected neighboring nodes which are not been masked. This term enforces high-order neighbors be selected based on their LCS values, reducing the influence of noisy or redundant neighbors.
To further mitigate the over-smoothing issue, we impose a mask constraint on high-order feature propagation:
(10) $\mathcal{L}_{\text{mask}}=\sum_{k=2}^{N}\|\tilde{\mathbf{A}}_{k}\|_{1},$
where $\tilde{\mathbf{A}}_{k}$ is the filtered high-order adjacency matrix, and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm. Combining all components, the final optimization objective is:
(11) $\mathcal{L}=\mathcal{L}_{\text{task}}+\lambda_{1}\mathcal{L}_{\text{LCS}}+\lambda_{2}\mathcal{L}_{\text{mask}},$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that control the balance between regularization and the primary task.
This optimization framework ensures that ScaleGNN maintains a balance between classification performance, robust high-order neighbor aggregation, and computational efficiency.
4.5. Time Complexity Analysis
@@
Let $n$ be the number of nodes, $m$ be the number of edges, and $f$ be the feature dimension, $K$ be the maximum number of hops. $a$ is the number of layers in MLP classifiers of SIGN, GAMLP, and GBP. $S$ in NARS denotes the number of subgraphs, $M$ in SeHGNN denotes the number of divided groups, and $R$ in RpHGNN is the number of node classes. In our proposed ScaleGNN, we perform the following major computations:
  • Multi-hop Adjacency Construction: Computing higher-order adjacency matrices requires $\mathcal{O}(Km)$, where $K$ is the maximum number of hops.
  • LCS-based Neighbor Filtering: Computing the Local Contribution Score (LCS) involves pairwise feature similarity for selected neighbors, yielding a complexity of $\mathcal{O}(K_{s}nf)$, where $K_{s}$ is the number of selected high-hop neighbors.
  • Feature Aggregation and Training: The GNN layer updates, including low-order and high-order feature aggregation, take $\mathcal{O}(nf^{2})$.
The final complexity per training iteration is approximately $\mathcal{O}(Km+K_{s}nf+nf^{2})$. Table 1 compares the time complexity of ScaleGNN with other scalable GNN architectures.
Table 1. Time complexity comparison of existing deep and scalable GNN models.
@@
S2GC $\mathcal{O}(Kmf)$ $\mathcal{O}(nf^{2})$ $\mathcal{O}(nf^{2})$
SIGN $\mathcal{O}(Kmf)$ $\mathcal{O}(anf^{2})$ $\mathcal{O}(anf^{2})$
GAMLP $\mathcal{O}(Kmf)$ $\mathcal{O}(anf^{2})$ $\mathcal{O}(anf^{2})$
GBP $\mathcal{O}(Kmf+K\sqrt{m}\log n)$ $\mathcal{O}(anf^{2})$ $\mathcal{O}(anf^{2})$
NARS $\mathcal{O}(KnSf+m)$ $\mathcal{O}(nf^{2})$ $\mathcal{O}(nf^{2})$
SeHGNN $\mathcal{O}(KnMf+m)$ $\mathcal{O}(nf^{2})$ $\mathcal{O}(nf^{2})$
RpHGNN $\mathcal{O}(KnRf+m)$ $\mathcal{O}(nf^{2})$ $\mathcal{O}(nf^{2})$
ScaleGNN $\mathcal{O}(Km+K_{s}nf)$ $\mathcal{O}(nf^{2})$ $\mathcal{O}(nf^{2})$
@@
Table 2. Statistical summaries of datasets.

### v4

@@ IV Methodology
In this section, we present ScaleGNN, an efficient end-to-end framework for scalable and robust GNN learning. ScaleGNN addresses over-smoothing, computational inefficiency, and redundant high-order information through three synergistic components: (1) Adaptive High-order Feature Fusion, (2) Low-order Enhanced Feature Aggregation, and (3) High-order Redundant Feature Masking. These modules enable ScaleGNN to process large-scale graphs efficiently while preserving essential information across neighborhood hops. The overall framework is illustrated in Fig. 1.
IV-A Adaptive High-order Feature Fusion
@@
Traditional GNNs are based on message passing through fixed adjacency matrices, where the adjacency structure remains constant throughout the layers. This mechanism can lead to over-smoothing, where node representations converge and lose their distinctiveness as the network depth increases. Furthermore, as nodes aggregate information from higher-order neighbors, the feature similarity across nodes can become excessive, thereby reducing the model’s expressiveness.
@@
To mitigate these issues, we propose a method for constructing distinct adjacency matrices for each neighborhood hop. This allows us to isolate the relationships specific to each hop, ensuring that each neighborhood level contributes uniquely to the node representation. We define the pure $i$-th hop adjacency matrix $\mathbf{A}_{i}$ as the difference between the $(i-1)$-th and $i$-th hops of the adjacency matrix:
$\mathbf{A}_{i}=\mathbf{A}^{i}-\mathbf{A}^{i-1},\quad i=1,2,3,\dots,K,$ (1)
@@
where $\mathbf{A}^{i}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is the $i$-th hop of the adjacency matrix, which captures the connections up to $i$ hops, and $\mathbf{A}_{i}$ isolates the new relations emerging specifically at the $i$-th hop. $K$ is the maximum number of hops, also referred to as the order. Specifically, both $\mathbf{A}_{1}$ and $\mathbf{A}^{1}$ are the adjacency matrix of the graph, and $\mathbf{A}^{0}$ is the zero matrix.
@@
Incorporating information from multiple hops can provide a more comprehensive view of the graph structure. However, the challenge lies in integrating these multiple hops in a way that preserves their distinct contributions. To address this, we introduce a learnable weight set $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\dots,\alpha_{K}]$, where each $\alpha_{i}\in\mathbb{R}$ ($i\in\{1,2,\ldots,K\}$) represents the specific important variant of the $i$-th hop adjacency matrix in the final aggregation:
$\tilde{\mathbf{A}}=\sum_{i=1}^{K}\alpha_{i}\mathbf{A}_{i},\quad\sum_{i=1}^{K}\alpha_{i}=1.$ (2)
@@
These weights are learned through back-propagation, ensuring that the model dynamically adjusts the contribution of each neighborhood hop. The use of the softmax function to optimize these weights ensures that contributions across different hops remain balanced. Finally, we obtain the adaptive high-order fused embedding matrix $\mathbf{H}\in\mathbb{R}^{|\mathcal{V}|\times d}$ as follows:
$\mathbf{H}=\sigma(\tilde{\mathbf{A}}\cdot\mathcal{X}\cdot\mathbf{W}),$ (3)
@@
where $d$ is the hidden embedding dimension, $\mathcal{X}\in\mathbb{R}^{|\mathcal{V}|\times f}$ is the input feature matrix, and $\mathbf{W}\in\mathbb{R}^{f\times d}$ is the learnable weight matrix for high-order features. The activation function $\sigma(\cdot)$ introduces non-linearity to the feature transformation.
@@
Specifically, we use the current model as our base version named ScaleGNNb, which is relatively simple and efficient.
IV-B Low-order Enhanced Feature Aggregation
@@
The integration of low-order (local) and high-order (global) features in many GNN-based models is done in a single and often static step. However, this can lead to high-order features dominating over low-order ones, which might diminish the model’s ability to preserve fine-grained local information. To address this, we propose explicitly separating low-order and high-order features before fusion. First, we obtain the low-order feature matrix $\mathbf{H}_{\text{low}}$ and the high-order feature matrix $\mathbf{H}_{\text{high}}$ as follows:
@@
$\mathbf{H}_{\text{low}}=\sigma(\mathbf{A}^{2}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{low}}),\quad\mathbf{H}_{\text{high}}=\sigma(\tilde{\mathbf{A}}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{high}}),$ (4)
@@
where $\mathbf{W}_{\text{low}},\mathbf{W}_{\text{high}}\in\mathbb{R}^{f\times d}$ are both learnable weight matrices for low-order and high-order features, respectively.
@@
To ensure a balanced integration of low-order and high-order features, we introduce a learnable balancing factor $\beta$, which allows for dynamic adjustment of diverse contributions from two feature types. The final node representation $\mathbf{H}$ is the weighted sum of low-order and high-order features:
$\mathbf{H}=\beta\cdot\mathbf{H}_{\text{low}}+(1-\beta)\cdot\mathbf{H}_{\text{high}},$ (5)
@@
where $\beta$ is optimized during training, ensuring the model can adaptively balance the trade-off between local and global information, depending on the specific task and graph structure.
@@
IV-C High-order Redundant Feature Masking
Higher-order neighbors introduce diverse, yet potentially redundant, information to GNNs. These neighbors will introduce noise without an effective masking mechanism, diluting the informative signal and impairing the model’s ability to generate useful node embeddings. To address this, we propose a novel method for selecting the most relevant and masking other redundant high-order neighbors based on their local contributions to node representations.
Inspired by the scaled dot-product attention mechanism [70], we introduce the Local Contribution Score (LCS) to assign importance scores to high-order neighbors. The LCS is designed to quantify the relevance of each neighboring node of the target node by considering both structural similarity and feature alignment. Specifically, for a node $v$ and its $i$-th hop neighbor $u$, the LCS value is computed as follows:
$\text{LCS}(v,u,i)=\frac{\exp\left(\frac{(\mathbf{W}_{1}\mathbf{x}_{v})^{\top}(\mathbf{W}_{2}\mathbf{x}_{u})}{\sqrt{d_{f}}}\right)}{\sum\limits_{u^{\prime}\in\mathcal{N}_{i}(v)}\exp\left(\frac{(\mathbf{W}_{1}\mathbf{x}_{v})^{\top}(\mathbf{W}_{2}\mathbf{x}_{u^{\prime}})}{\sqrt{d_{f}}}\right)}.$ (6)
Here, $\mathbf{x}_{v},\mathbf{x}_{u}\in\mathbb{R}^{f}$ are the input feature vectors of the target node $v$ and its $k$-hop neighbor $u$. $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{d_{f}\times f}$ are learnable projection matrices that map features into a latent attention space of dimension $d_{f}$, which is a tunable hyperparameter, and typically $d_{f}\leq f$ to ensure efficiency. $\mathcal{N}_{i}(v)$ denotes the $i$-hop neighborhood of node $v$ induced by $\mathbf{A}_{i}$. The denominator normalizes attention scores within each local $i$-hop neighborhood using the softmax function. This equation allows the model to learn context-aware neighbor importance scores in a differentiable and robust way.
The LCS allows us to mask redundant high-order neighbors and identify the most relevant neighbors by considering both local structure and feature similarity, enabling the model to selectively pass information from the most meaningful high-order neighbors. Once the LCS is computed for each neighbor, we select the top-$m_{i}$ neighbors with the highest LCS values for each node among its $i$-hop neighbors:
$\mathbf{A}_{i}^{\text{filter}}=\text{Top}_{m_{i}}(\mathbf{A}_{i},\text{LCS}),$ (7)
[…]
IV-D Joint Optimization
To ensure the effectiveness of our proposed ScaleGNN, we design an optimization framework that consists of three key loss components: the node classification task loss, the LCS regularization loss, and the sparse constraint loss on the high-order information. These losses work together to enhance model performance, suppress noise from uninformative high-order neighbors, and maintain computational efficiency.
@@ IV-D Joint Optimization
$\mathcal{L}_{Cla}=-\sum_{i\in\mathcal{V}}y_{i}\log\hat{y}_{i},$ (8)
where $y_{i}$ represents the one-hot vector of the ground-truth class label for node $i$, and $\hat{y}_{i}$ is the predicted probability distribution over classes generated by a softmax layer.
To further mitigate the over-smoothing issue, we impose a sparse constraint (SC) on high-order feature propagation:
$\mathcal{L}_{SC}=\sum_{i=2}^{K}\|\bar{\mathbf{A}}_{i}^{\text{filter}}\|_{1},$ (9)
where $\bar{\mathbf{A}}_{i}^{\text{filter}}$ is the filtered high-order adjacency matrix, $\bar{\mathbf{A}}_{i}^{\text{filter}}=\mathbb{I}\left[\mathbf{A}_{i}^{\text{filter}}>0\right]$, where $\mathbb{I}[\cdot]$ denotes the element-wise indicator function, and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm. This term constrains the model to minimize the number of neighbors retained for each node at each hop.
To guide the model in selecting meaningful high-order neighbors, we also introduce an LCS-based regularization:
$\mathcal{L}_{\text{LCS}}=\sum_{v\in\mathcal{V}}\sum_{i=2}^{K}\sum_{u\in\mathcal{N}_{i}^{\text{select}}(v)}\left(1-\text{LCS}(v,u,i)\right)^{2},$ (10)
where $\mathcal{N}_{i}^{\text{select}}(v)$ is the retained neighboring nodes which are not been filtered. This term enforces high-order neighbors to be selected based on their LCS values, reducing the influence of noisy or redundant neighbors.
Combining all components, the final optimization objective is as follows:
$\mathcal{L}=\mathcal{L}_{Cla}+\lambda_{1}\mathcal{L}_{\text{LCS}}+\lambda_{2}\mathcal{L}_{SC},$ (11)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters balancing regularization against the primary task objective. Specifically, enforcing sparsity constraints during the selection of higher-order neighbors per hop allows the model to strike a balance between preserving critical features and mitigating redundant information.
Overall, this optimization framework ensures that our ScaleGNN maintains a balance between node classification performance, robust high-order neighbor aggregation, and computational efficiency.
Algorithm 1 ScaleGNN
0:  Graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X})$, adjacency $\mathbf{A}$, max order $K$
0:  Node representations $\mathbf{H}$
1:  Construct pure $i$-hop adjacency:
2:  for $i=1$ to $K$ do
3:     $\mathbf{A}_{i}=\mathbf{A}^{i}-\mathbf{A}^{i-1}$
4:  end for
5:  Adaptive high-order feature fusion:
6:  Learn weights $\alpha_{1},\dots,\alpha_{K}$ via softmax, $\sum_{i}\alpha_{i}=1$
7:  $\tilde{\mathbf{A}}=\sum_{i=1}^{K}\alpha_{i}\mathbf{A}_{i}$
8:  $\mathbf{H}_{\text{high}}=\sigma(\tilde{\mathbf{A}}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{high}})$
9:  Low-order feature aggregation:
10:  $\mathbf{H}_{\text{low}}=\sigma(\mathbf{A}^{2}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{low}})$
11:  High-order redundant feature masking:
12:  for each node $v$ do
13:     for $i=2$ to $K$ do
14:        for each neighbor $u$ in $\mathbf{A}_{i}$ do
15:           Calculate $\text{LCS}(v,u,i)$ by Eq. (6)
16:        end for
17:        Select top-$m_{i}$ neighbors by LCS, mask others in $\mathbf{A}_{i}$
18:     end for
19:  end for
20:  Update $\tilde{\mathbf{A}}$ and $\mathbf{H}_{\text{high}}$ using $\bar{\mathbf{A}}_{i}^{\text{filter}}$
21:  Feature fusion:
22:  Learn $\beta\in[0,1]$
23:  $\mathbf{H}=\beta\mathbf{H}_{\text{low}}+(1-\beta)\mathbf{H}_{\text{high}}$
24:  return  $\mathbf{H}$
Algorithm 1 summarizes the overall process of the proposed ScaleGNN framework. The model first constructs pure $k$-hop adjacency matrices to isolate neighbor information at different topological scales. It then performs adaptive high-order feature fusion by learning attention weights across multiple hops, and separately aggregates low-order features via the original adjacency. To suppress redundant information introduced by distant neighbors, a local contribution score (LCS) is used to select the most informative nodes, refining the high-order structure. Finally, the node representations are obtained by weighted fusion of the low- and high-order embeddings.
IV-E Time Complexity Analysis
@@
Let $|\mathcal{V}|$ be the number of nodes, $|\mathcal{E}|$ be the number of edges, and $f$ be the feature dimension, $K$ be the maximum number of hops, $n$ be the maximum number of top-$m_{i}$ ($i\in\{1,2,\ldots,K\}$) neighbors. $L$ is the number of layers in MLP classifiers of SIGN, GAMLP, and GBP, $S$ in NARS denotes the number of subgraphs, $P$ in SeHGNN denotes the number of divided groups, and $R$ in RpHGNN is the number of node classes. In our proposed ScaleGNN, we perform the following major computations:
  • Multi-hop Adjacency Construction: To construct order-distilled adjacency matrices $\{\mathbf{A}_{i}\}_{i=1}^{K}$ where $\mathbf{A}_{i}=\mathbf{A}^{i}-\mathbf{A}^{i-1}$, we perform $K$ sparse matrix multiplications. The cost per multiplication is $O(|\mathcal{E}|)$ in sparse setting, so the total cost is $O(K|\mathcal{E}|)$.
  • LCS-based Neighbor Filtering: For each node $v$ and $k$-hop neighbor $u$, attention-based LCS involves two linear projections ($\mathbf{W}_{1}\mathbf{x}_{v}$, $\mathbf{W}_{2}\mathbf{x}_{u}$), a dot-product, and softmax computation: (1) Projection cost: amortized $O(|\mathcal{V}|fd_{f})$ (precomputed). (2) Attention score computation: $O(K|\mathcal{V}|nd_{f})$. (3) top-$m_{k}$ filtering: $O(K|\mathcal{V}|n\log n)$ using efficient heap sort. The overall complexity is approximately $O(Kn|\mathcal{V}|f)$ and remains linear in retained neighbors and supports batch-wise parallel execution.
  • Feature Aggregation and Training: Once filtered adjacency matrices are obtained, the time complexity of aggregation step is $O(Kn|\mathcal{V}|f)$. The time complexity of training and inference is $O(|\mathcal{V}|f^{2})$
Finally the end-to-end time complexity of our ScaleGNN is approximately $O(K|\mathcal{E}|+Kn|\mathcal{V}|f+|\mathcal{V}|f^{2}$).
Table I compares the time complexity of ScaleGNN with other scalable GNN architectures.
TABLE I: Time complexity comparison of existing deep and scalable GNN models.
@@
S2GC $\mathcal{O}(K|\mathcal{E}|f)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
SIGN $\mathcal{O}(K|\mathcal{E}|f)$ $\mathcal{O}(L|\mathcal{V}|f^{2})$ $\mathcal{O}(L|\mathcal{V}|f^{2})$
GAMLP $\mathcal{O}(K|\mathcal{E}|f)$ $\mathcal{O}(L|\mathcal{V}|f^{2})$ $\mathcal{O}(L|\mathcal{V}|f^{2})$
GBP $\mathcal{O}(K|\mathcal{E}|f+K\sqrt{|\mathcal{E}|}\log|\mathcal{V}|)$ $\mathcal{O}(L|\mathcal{V}|f^{2})$ $\mathcal{O}(L|\mathcal{V}|f^{2})$
NARS $\mathcal{O}(K|\mathcal{V}|Sf+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
SeHGNN $\mathcal{O}(K|\mathcal{V}|Pf+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
RpHGNN $\mathcal{O}(K|\mathcal{V}|Rf+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
ScaleGNN $\mathcal{O}(Kn|\mathcal{V}|f)+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
@@
TABLE II: Statistical summaries of datasets.